OpenStudy (anonymous):

You are driving along a country road when you suddenly notice a log in the road ahead of you and immediately apply your brakes. You travel a distance of s(t) = 66 t - 6 t^2 feet in t seconds after you hit the brakes. The log was 180 feet ahead of you when you noticed it, which unfortunately wasn't soon enough. At what time (in seconds) did you hit the log? How fast were you going (in ft/sec) when you hit it?

OpenStudy (anonymous):

which grade is it?

OpenStudy (anonymous):

well since 180 is the distance you know distance is a function of time. so if you set the function = to 180 then you can solve for t

OpenStudy (anonymous):

then your speed will be distance/time.

OpenStudy (anonymous):

oh i see i need a s = d/t equation

OpenStudy (anonymous):

dude theres something wrong with the equation it is coming 2094ft/sec

OpenStudy (anonymous):

what did you get for your time.

OpenStudy (anonymous):

well u differentiate the equation and then put 180 in place of t and u get ur speed

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

180 is not a time.

OpenStudy (anonymous):

so i got s(t) = 66-6t^2 v(t) = -12t a(t) = -12

OpenStudy (anonymous):

what else will u substitute for t dude

OpenStudy (anonymous):

s(t)=180 because that is distance as a function of time.

OpenStudy (anonymous):

yeah jgeurts u got it right

OpenStudy (anonymous):

if s(t) = 180 then i get t^2 = -19???

OpenStudy (anonymous):

dont do it that way what u got is right jgeurts

OpenStudy (anonymous):

looks like you dropped a t when you took the derivative of your first function it should have been 66-12t

OpenStudy (anonymous):

as long as the first function wasnt a typo. because i see 66t-6t^2=s(t)

OpenStudy (anonymous):

what are u saying man s(t)=180 has t^2 and not 12 t

OpenStudy (anonymous):

so how fast was i going?

OpenStudy (anonymous):

i was just correcting the first derivative of the original function.

OpenStudy (anonymous):

-12t

OpenStudy (anonymous):

but i can't solve for t

OpenStudy (anonymous):

u don't have to u can write the speed in terms of a variable

OpenStudy (anonymous):

i need to

OpenStudy (anonymous):

was your original function s(t)=66t-6t^2?

OpenStudy (anonymous):

yes….

OpenStudy (anonymous):

then your first derivative is wrong. it should have a constant in it.

OpenStudy (anonymous):

so i got s(t) = 66t-6t^2 v(t) = 66-12t a(t) = -12

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

yeah and t=5 or 6 i mean it has 2 values

OpenStudy (anonymous):

yep. but you still have to find your time. so the only way i can see to do that is to use your distance on your original function to get time.

OpenStudy (anonymous):

dude i just gave u the value of time

OpenStudy (anonymous):

how do i do distance on my original function?

OpenStudy (anonymous):

thanks Rahul, but how did u get it?

OpenStudy (anonymous):

your distance is 180 feet from the log. and dude.. i posted that at the same time as you. get your panties out of a bunch...

OpenStudy (anonymous):

180=66t-6t^2 like i said a while back.

OpenStudy (anonymous):

u solve the eqn 66t-6t^2=180 get t then put it in -12t to get the speed

OpenStudy (anonymous):

are we done

OpenStudy (anonymous):

-12t is not the first derivative of the position function here.

OpenStudy (anonymous):

66-12t is

OpenStudy (anonymous):

sorry i mean 66-12t

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

hey don't i get a medal?

OpenStudy (anonymous):

66t -6t^2 = 180 t = sqrt19

OpenStudy (anonymous):

plug sqrt19 into v(t)?

OpenStudy (anonymous):

dude jgeurts do u even know maths

OpenStudy (anonymous):

the problem isnt finished. uh. thats not what i was thinking it should be 5 or 6. hey man. there is no reason to be rude.

OpenStudy (anonymous):

lol im so bad, I'm going off that original screwed up equation

OpenStudy (anonymous):

your 2 possible answers for time are 5 and 6 since its a quadratic equation you will have 2 possible answers.

OpenStudy (anonymous):

it is 5 or 6 help urself

OpenStudy (anonymous):

you will take your time and plug it into your velocity function to get your velocity.

OpenStudy (anonymous):

ok, its 5 then plug it into 66 -12t so that would be 6s?

OpenStudy (anonymous):

100% thank you both!

OpenStudy (anonymous):

6s? what do you mean by that. you will have 6 as your answer. do you mean 6 feet per second?

OpenStudy (anonymous):

yeah fps

OpenStudy (anonymous):

sweet. then that is correct for time 5